OpenStudy (princesssleelee):

Given the trinomial 2x2 + 4x + 4, predict the type of solutions. Two rational solutions One rational solution Two irrational solutions Two complex solutions

OpenStudy (princesssleelee):

i found out the solution is 2(x2+2x+2)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle x^2+2x+1=(x+1)^2 }\) (Right ?) \(\color{#000000 }{ \displaystyle 2x^2+4x+4=0 }\) \(\color{#000000 }{ \displaystyle x^2+2x+2=0 }\) (divided each term by 2) \(\color{#000000 }{ \displaystyle (x^2+2x+1)+1=0 }\) \(\color{#000000 }{ \displaystyle (x^2+2x+1)=-1 }\)

OpenStudy (solomonzelman):

What I am showing is called "completing the square [method]". If you need an example of that method, let me know....

OpenStudy (princesssleelee):

But im not sure of the answer choices, i was thinking "D" however it concludes in ONE solution

OpenStudy (solomonzelman):

I wanted you to find the two solutions and then for you to choose the answer.... I will give you a hint for the answer after/by proceeding \(\color{#000000 }{ \displaystyle x^2+2x+1=-1 }\) \(\color{#000000 }{ \displaystyle (x+1)^2=-1 }\) \(\color{#000000 }{ \displaystyle \sqrt{(x+1)^2}=\sqrt{-1} }\)

OpenStudy (solomonzelman):

from there you should tell me what kind of answers/solutions you will get

OpenStudy (princesssleelee):

im not sure what each one means

OpenStudy (princesssleelee):

Im stuck between D and C though, the other ones dont seem logical to me

OpenStudy (solomonzelman):

irrational number is [the result from] any Nth root of a non-perfect Nth. (short but perhaps abstruse definition) Also, irrational numbers include \(\pi\), \(e\), and others such...

OpenStudy (princesssleelee):

So it eliminates answer choice A and B

OpenStudy (solomonzelman):

Complex = Imaginary Irrational = Real numbers that are 'approximate'

OpenStudy (princesssleelee):

Leaving answer choice D, because -1 is equal to "i" (imaginary)

OpenStudy (triciaal):